OpenStudy (anonymous):

What is the solution set of 6x 2 - 24 = 0

jimthompson5910 (jim_thompson5910):

First add 24 to both sides \[\Large 6x^2 - 24 = 0\] \[\Large 6x^2 = 24\]

jimthompson5910 (jim_thompson5910):

Then divide both sides by 6 \[\Large 6x^2 = 24\] \[\Large x^2 = \frac{24}{6}\] \[\Large x^2 = 4\]

jimthompson5910 (jim_thompson5910):

From here, how do we solve \[\Large x^2 = 4\] for x?